OpenStudy (anonymous):

The speed of a passenger train is 14mph faster than the speed of a freight train. The passanger train travels 290 miles in the same time it takes the freight train to travle 220 miles. Find the speed of both trains........ would I use r=d/t?

OpenStudy (anonymous):

Sort of. v=d/t t=d/v d1/v=d2/v2 (220)/v=(290)/(v+14) v in this case is the speed of the freight train. solve for v

OpenStudy (anonymous):

k thanks :)